OpenStudy (lgbasallote):

The ratio of helper to learner in OpenStudy is 1:20 during regular time. Over the summer, the total population of OpenStudy decreased by 32%. What will be the ratio of helper to learner in the new population?

OpenStudy (anonymous):

haha got to love this question! 1:12

OpenStudy (lgbasallote):

how?

OpenStudy (saifoo.khan):

I think we will find the 32% of 21 first.

OpenStudy (saifoo.khan):

i mean decrease 32 %/

OpenStudy (lgbasallote):

21 is the totalright?

OpenStudy (lgbasallote):

total ratio i mean

OpenStudy (saifoo.khan):

No. 20+1 = 21 So Yes. lol.

OpenStudy (saifoo.khan):

6.72 decreased.

OpenStudy (saifoo.khan):

21 - 6.72 = 14.28

OpenStudy (saifoo.khan):

idk how to continue from here.

OpenStudy (lgbasallote):

hah lol

OpenStudy (saifoo.khan):

@dpaInc , try this out!

OpenStudy (saifoo.khan):

im not sure but i will try it out, 1/21 * 100 = 4.76% were the helpers. ON TOTAL. And since the total popl decreased, the ratio should remain the same (i think) Next, thinking..

OpenStudy (anonymous):

thing is... is that 32% include the helpers? when u say a decrease total of 32%, i'm assuming the whole population so shouldn't the ratio still be 1:20? idk... :(

OpenStudy (lgbasallote):

so it;s still same?

OpenStudy (anonymous):

is the 32% decrease only the ones being helped? if yes, then it wouldn't be the same...

OpenStudy (saifoo.khan):

@lgbasallote , so you asked a wrong question! D:

OpenStudy (saifoo.khan):

@dpaInc 's way seems legit.

OpenStudy (lgbasallote):

lol i just copied it from my book :P blame the book

OpenStudy (anonymous):

umm don't bank on my answer... i only puking out what i think...:)

OpenStudy (saifoo.khan):

Books are waste of trees. Dont buy them.

OpenStudy (anonymous):

Ratio will remain same even after reduction in total population , assuming reduction is uniform across the population.

OpenStudy (anonymous):

what' happens when we get a an electromagnetic pulse from the nuclear warheads? all those hard drives will be wiped clean... so books are good...:)

OpenStudy (anonymous):

The ratio will be unchanged!

OpenStudy (anonymous):

watch the move "Book of Eli".... good story...

OpenStudy (saifoo.khan):

If others want to see the trailer here it is: http://www.youtube.com/watch?v=kAMUv22y1og

OpenStudy (saifoo.khan):

Nukes are cool. B) It makes you look cooler. ;) Hard drives still aren't bad for trees. lol.

OpenStudy (anonymous):

hmm... we always seem to go off on tangents... don't you think?

OpenStudy (anonymous):

good point..

OpenStudy (saifoo.khan):

Lol. i like going off the tangents. :D Don't you? I simply love it.

OpenStudy (anonymous):

agree...:)

OpenStudy (saifoo.khan):

;)